The density of He is 0.1784kg/m^3 at STP . If a given mass of He at STP is allowed to expand to 1.5 times its initial volume by changing the temperature and pressure, compute its resultant density.(Ans:-0.1189kg/m^3)

Grade:11

3 Answers

Vikas TU
7 years ago
From charles law moles are to be constant.
P1V1/T1 = P2V2/T2
1*V/273 = P2*1.5V/T2
P2/T2 = 1/(273 x 1.5)..............(1)
 
New density = P2M/RT2 = >  (P2/T2)*(M/R) = > 1/(273 x 1.5) * (4/0.0821) = > 0.1189 kg/m^3
Vaibhavi
7 years ago
How the the No of moles are constant plz explain?
Madhumit
5 years ago
The density of he is 0.178 4 kg per metre cube at STP if a given mass of he at STP is allowed to expand to 1.5 times its initial volume by changing the temperature and pressure compute its resultant density